OpenStudy (anonymous):

[find f ^{-1}(x) f(x)=\sqrt[3]{x-7}\]

jimthompson5910 (jim_thompson5910):

Replace f(x) with y. Swap x and y. Then solve for y to get the inverse. \[\Large f(x)=\sqrt[3]{x-7}\] \[\Large y=\sqrt[3]{x-7}\] \[\Large x=\sqrt[3]{y-7}\] \[\Large x^3=y-7\] \[\Large x^3+7=y\] \[\Large y=x^3+7\] \[\Large f^{-1}(x)=x^3+7\]

OpenStudy (anonymous):

though I'm not quite sure how you solved it, thanks for your help, I will try these steps

jimthompson5910 (jim_thompson5910):

you undo each operation so you first undo the cube root by cubing both sides then you undo the -7 by adding 7 to both sides